OpenStudy (anonymous):

Please help! I will give a medal if you can help me!!! Donna teaches dance in groups of five students. She rents a hall for the dance classes at $450 per month. She charges $40 per hour per student. Donna teaches four groups of students, and her net earnings per month are $4,350. Part A: If each student learns dance for the same number of hours, write an equation that can be used to calculate the number of hours each student learns dance in a month. (5 points) Part B: How many hours did each student learn dance each month? Show your work and justify each step. (5 points)

OpenStudy (anonymous):

Wait i am reading the question.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Part A student × hours × 40= 4350+450=4800 +450 because even when she pays 350 she gets 4350.

OpenStudy (anonymous):

hours = 4800 ÷ 40 = 120 120 ÷ 5 = 24 because there are 5 students.

OpenStudy (anonymous):

Okay..... that makes sense

OpenStudy (anonymous):

Thanks!!!

OpenStudy (anonymous):

yw :)

OpenStudy (anonymous):

you are free to ask any questions if you didn't get something

OpenStudy (anonymous):

No it makes more sense now it's just when I read it the first time I was really confused and I am tired from doing a bunch of Algebra.

OpenStudy (anonymous):

@JCMMusic did you get this right ?